# How do Pre-Trained Models Support Software Engineering? An Empirical Study in Hugging Face

### v1

@@ V-B Updatability
The replication package [21] accompanying this work includes the complete pipeline described in Section III, enabling full replicability of our results. By following the documented steps, users can reproduce the same classification of PTMs available up to March 24, 2025, while also retrieving and classifying PTMs added since that date. This supports replicability and allows others to validate or extend our work.
However, given the growing number of PTMs in HF [60, 8], it is not necessary to rerun the full pipeline unless modifications are required (e.g., updating the taxonomy of SE tasks or validating the classification strategy). To keep the dataset up to date, users can retrieve only the new PTMs added since our cutoff date and apply the pipeline exclusively to those. The resulting classified PTMs can then be merged with our dataset, allowing efficient updatability.
VI Threats to Validity
@@
Potential threats impacting our study’s validity are outlined.
@@
External Validity: While HF is a leading platform for sharing PTMs, other repositories, such as PyTorch Hub [61], may host relevant PTMs that are not captured in our analysis. Moreover, ongoing changes in HF’s platform structure, documentation practices, or model metadata could affect the broader applicability of our conclusions.
Construct Validity: Our classification of SE tasks and activities builds upon an established taxonomy derived from existing literature [16]. We extended this taxonomy to include a broader set of 147 SE tasks, aiming for more granular coverage. Nevertheless, it may evolve as the SE community introduces new standards or refinements, which could impact future interpretations or mappings.
Conclusion Validity: We acknowledge the use of Gemini 2.0 Flash to support part of the classification process. To mitigate potential risks, we designed a pilot evaluation protocol with five distinct prompting scenarios (one per SE activity), and validated outputs on a statistically significant sample. A total of 1,346 samples were manually reviewed across all five pilots, involving three independent annotators. This process enabled us to assess the consistency, accuracy, and potential biases in the model’s responses.
@@
VII Conclusion and Future Work
@@
For future work, we plan to extend this classification to other repositories, such as Papers with Code [62], PyTorch Hub [61], and TensorFlow Hub [63], allowing for a broader analysis of PTMs in SE. We also aim to enhance the classification by recovering and integrating relevant information from original base model cards, which may not be explicitly stated in derivative model cards. To make our classification more actionable, we propose developing a dashboard that enables automatic sampling and selection of PTMs for real-world applications, supporting both researchers in conducting empirical studies and practitioners in integrating PTMs into software engineering workflows. Finally, we intend to explore how this classification can be embedded into ASE pipelines by enabling automatic recommendation and adaptation of PTMs to SE tasks, such as design pattern or bug detection.
VIII Artifact availability
To support transparency, reproducibility, and reuse, we provide a publicly available replication package containing all research artifacts used in this study [21]. It includes the taxonomy of 147 software engineering tasks aligned with the SDLC, along with the raw and preprocessed data, and the full pipeline for data collection, preparation, search, and filtering. We also share the manual annotations made by human experts during the study.

### v2

@@ 5.5 Updatability
The replication package replication_package accompanying this work includes the complete pipeline described in Section 3, enabling full replicability of our results. By following the documented steps, users can reproduce the same cataloguing of PTMs available up to March 24, 2025, while also retrieving and classifying PTMs added since that date. This supports replicability and allows others to validate or extend our work.
Given the exponential growth of PTMs in HF ait2025suitability; 10.1145/3643991.3644898, re-executing the full pipeline for the entire repository is computationally inefficient and arguably unnecessary unless the taxonomy itself requires modification (e.g., updating the taxonomy of SE tasks or validating the classification strategy). To maintain an up-to-date dataset, our framework supports incremental updates: users can retrieve only the PTMs created after the cutoff date, catalogue them, and merge the results with our baseline dataset. This ensures efficient maintainability without redundant processing.
We acknowledge that a research paper is inherently static (our analysis captures the state of HF as of March 24, 2025), whereas the HF ecosystem is highly dynamic. Since our data collection, the number of repositories has continued to grow at a rapid pace laufer2025anatomy. Given that hundreds of new repositories are created daily  [^3] https://huggingface.co/posts/clem/238420842235482, the absolute numbers presented in this article will inevitably become outdated (as in any other article of this kind). To address this temporal limitation, the value of this work lies not only in the static snapshot provided but in the updatability of the framework. While the raw numbers presented in this article reflect the state of HF as of March 2025, we argue that the distribution of SE tasks and the utility of our taxonomy remain robust. By applying our validated filtering and cataloguing methodology to new entries, the community can maintain a living catalogue. This ensures that the insights regarding the distribution of SE tasks and the utility of the taxonomy remain relevant, even as the underlying raw numbers of the HF ecosystem continue to rise.
Furthermore, our catalogue is designed to be extensible and interoperable with other existing datasets. Although our work focuses specifically on SE, it can be integrated with broader, generic catalogues such as HFCommunity 10123660 or PeaTMOSS jiang2024peatmoss.
6 Threats to Validity
@@
Potential threats that could impact our study’s validity are outlined.
@@
  • External Validity: While HF is a leading platform for sharing PTMs, other registries, such as PyTorch Hub pytorchPyTorch, may host relevant PTMs that are not captured in our analysis. Moreover, ongoing changes in HF’s platform structure, documentation practices, or model metadata could affect the broader applicability of our conclusions.
  • Construct Validity: Our catalogue of SE tasks and activities builds upon an established taxonomy derived from existing literature 10.1145/3695988. We extended this taxonomy to include a broader set of 147 SE tasks, aiming for more granular coverage. Nevertheless, it may evolve as the SE community introduces new standards or refinements, which could impact future interpretations or mappings.
  • Conclusion Validity: While manual annotation may introduce subjective bias, we mitigated this threat through inter-coder agreement practices. We acknowledge the use of Gemini 2.0 Flash to support part of the classification process. To mitigate potential risks, we designed a pilot evaluation protocol with five distinct prompting scenarios (one per SE activity), and validated outputs on a statistically significant sample. A total of 1,346 samples were manually reviewed across all five pilots, involving three independent annotators. This process enabled us to assess the consistency, accuracy, and potential biases in the model’s responses.
@@
7 Conclusion and Future Work
@@
For future work, we plan to extend this classification to other registries, such as GitHub Models githubGitHubModels and PyTorch Hub pytorchPyTorch, allowing for a broader analysis of SE PTMs. We also aim to enhance the catalogue by recovering and integrating relevant information from original base model cards, which may not be explicitly stated in derivative model cards. Finally, we intend to explore how this classification can be embedded into SE pipelines by enabling automatic recommendation and adaptation of PTMs to SE tasks, such as design pattern or bug detection.
Statements and Declarations
Funding
This work was supported by Grant PID2024-156019OB-I00 funded by MICIU/AEI/10.13039/501100011033 and by ERDF, EU. Alexandra González additionally thanks the FI-STEP grant 2025 STEP-00407.
Ethical approval
This study does not involve human participants or animals.
Informed consent
No human subjects were involved in this study.
Author Contributions
  • Alexandra González: Data Collection, Data Processing, Data Validation, Data Analysis, Writing - Original Draft.
  • Xavier Franch: Supervision, Data Validation, Writing - Review & Editing
  • David Lo: Writing - Review & Editing
  • Silverio Martínez-Fernández: Supervision, Data Validation, Writing - Review & Editing
Data Availability Statement
To support transparency, reproducibility, and replicability, and in line with the EMSE open-science guidelines  [^4] https://github.com/emsejournal/openscience, we provide a publicly available replication package containing all research artifacts used in this study replication_package. It includes the taxonomy of 147 software engineering tasks aligned with the SDLC, along with raw and preprocessed data, and the complete pipeline for data collection, preparation, search, and filtering. We also share the manual annotations made by human experts during the study.
Conflict of Interest
The authors declare that they have no conflict of interest.
Clinical Trial Number
Not applicable.